Mathematics 241 – Multivariable Calculus Practice 1 – August 31, 2018

Your Name:

A) (10) Sketch or describe in words the set C = {(x, y) ∈ R^{2} : (x − 1)^{2} + (y − 1)^{2} = 4}.

B) (10) Give a set notation description for the plane in R^{3} parallel to the yz-plane and containing the point (−1, 2, 3).